# Optimizing post-Newtonian parameters and fixing the BMS frame for numerical-relativity waveform hybridizations

### v1

@@ Abstract
Numerical relativity (NR) simulations of binary black holes provide precise waveforms, but are typically too computationally expensive to produce waveforms with enough orbits to cover the whole frequency band of gravitational-wave observatories. Accordingly, it is important to be able to hybridize NR waveforms with analytic, post-Newtonian (PN) waveforms, which are accurate during the early inspiral phase. We show that to build such hybrids, it is crucial to both fix the Bondi-Metzner-Sachs (BMS) frame of the NR waveforms to match that of PN theory, and optimize over the PN parameters. We test such a hybridization procedure including all spin-weighted spherical harmonic modes with $|m|\leq\ell$ for $\ell\leq 8$, using 29 NR waveforms with mass ratios $q\leq 10$ and spin magnitudes $|\chi_{1}|,|\chi_{2}|\leq 0.8$. We find that for spin-aligned systems, the PN and NR waveforms agree very well. The difference is limited by the small nonzero orbital eccentricity of the NR waveforms, or equivalently by the lack of eccentric terms in the PN waveforms. To maintain full accuracy of the simulations, the matching window for spin-aligned systems should be at least 5 orbits long and end at least 15 orbits before merger. For precessing systems, the errors are larger than for spin-aligned cases. The errors are likely limited by the absence of precession-related spin-spin PN terms. Using $10^{5}\,M$ long NR waveforms, we find that there is no optimal choice of the matching window within this time span, because the hybridization result for precessing cases is always better if using earlier or longer matching windows. We provide the mean orbital frequency of the smallest acceptable matching window as a function of the target error between the PN and NR waveforms and the black hole spins.
@@ I Introduction
Such hybrid models were developed for example, in Refs. [25, 26, 27, 28, 29, 30]. These models attained mismatches of $10^{-4}$ between the hybrid and NR waveforms in the matching window (for $m\neq 0$ modes) for non-precessing systems [29], and between $10^{-2}$ and $10^{-1}$ (for $m\neq 0$ modes) for precessing systems [30]. However, these models have four major limitations: (1) They did not account for the $m=0$ modes since their NR waveforms did not include the correct gravitational-wave memory contribution. (2) The NR waveforms used in these models were not in the same frame as the PN waveforms. Reference [31] showed that comparing CCE waveforms without fixing the gauge freedom can result in relative errors three orders of magnitude larger than if one first fixes such freedoms. And they also showed that comparing CCE waveforms after mapping them to the same BMS frame can achieve smaller errors than comparing the extrapolated waveforms. Furthermore, Ref. [31] successfully developed a hybrid model that incorporates memory-containing waveforms. (3) These models used NR values of masses and spins to compute the corresponding PN waveform. As we will demonstrate in this paper, the NR values of these parameters differ from the PN values that minimize mismatch. (4) Existing models are primarily applicable to spin-aligned or anti-aligned systems, such as binaries born from isolated stellar evolution [32, 33, 34]. The accuracy of hybridization for precessing systems is severely limited by the inaccurate methodology previously mentioned: the mismatch is on the order of $10^{-2}$ or higher [30]. Theoretical considerations [32, 34] and observational evidence [9, 35, 36] indicate that spins can be randomly oriented for binaries that form dynamically in dense environments, and even binaries formed in isolation can exhibit distinctive precessional dynamics. Therefore, it is crucial to develop accurate hybrid models that are valid for precessing systems.
@@
One technique that enables the extraction of waveforms at null infinity is Cauchy-characteristic evolution (CCE) [37, 38, 39, 40, 41]. It uses a worldtube from the Cauchy evolution of the Einstein field equations as the inner boundary to a characteristic evolution along null rays. The gravitational information is propagated and evolved to future null infinity, so one can obtain waveforms in any desired gauge. Reference [31] has shown that by mapping the CCE waveform to the correct Bondi-Metzner-Sachs (BMS) frame, the NR waveform agrees much better with that of PN. A key advantage of the hybrid waveforms presented here is that we use CCE NR waveforms. By contrast, previous hybrid waveform models [25, 26, 27, 28, 29, 30] used NR waveforms that were computed at null infinity using an extrapolation technique [42] that does not have a well-defined BMS frame and does not yield the correct memory contribution.
@@
In this work, we simultaneously optimize over the PN parameters and fix the BMS frame of NR waveforms to build hybrid waveforms. The optimization and the fixing of the BMS frame, which includes the alignment of the rotation and time shift, are performed inside a time interval called a matching window. Because we use CCE and thus obtain the correct gravitational-wave memory, we are able to include all $-\ell\leq m\leq\ell$ for $\ell\leq 8$ spin-weighted spherical harmonic modes in the hybrid waveforms. We test the hybridization procedure using 29 NR waveforms from the SXS catalog [43] with mass ratios $q\leq 10$ and $|\chi_{1}|,|\chi_{2}|\leq 0.8$. The length of these waveforms ranges from 60 to over 100 orbits, making it possible to test the hybrid waveforms by comparing them to long NR waveforms at a very early stage. We find that for spin-aligned systems, we can reduce the PN–NR mismatches to the error caused by small non-zero eccentricities of the NR systems. We achieve a mismatch (in a 10-orbit long matching window ending $8000\,M$ before merger) around two to three orders of magnitude smaller than previous models [25, 26, 27, 28, 29]. For precessing systems, we find that hybridization is likely limited by the absence of precession-related spin-spin PN terms, yielding mismatches one to two orders of magnitude smaller than previous models [44].
@@
This paper is organized as follows. In Sec. II, we introduce the PN and NR waveforms that we use to build the hybrid waveforms, and we discuss the freedoms in the physical parameters, 3D rotations and time shift, and further BMS frame parameters not previously considered. In Sec. III, we explain how we match the physical parameters and inertial frames between NR and PN waveforms. Following this, we present the complete procedure to build the hybrid waveforms in Sec. IV. In Sec. V, we show the hybridization results as well as the improvement from PN parameters and BMS frame fixing compared with previous models, and we use the findings to more precisely compare the PN and NR waveforms. Sec. VI analyzes the sources of error for spin-aligned and precessing systems. Additionally, based on our error analysis, we make recommendations for future surrogate models’ matching window selection in Sec. VII. In Sec. VIII, we provide a few closing remarks. We make the code used for this hybridization procedure publicly available in Ref. [45].
@@ II.1 Post-Newtonian waveforms
In this study, we use PN theory to describe the GW waveforms of BBHs in quasi-circular orbits during the early inspiral stage, and include GW memory contributions. Memory effects should be inherent in any correct PN formulation, but their importance can often be overlooked. In our case, we find that accounting for memory is crucial when comparing to NR waveforms that have memory, and thus crucial for accurate hybridization.
@@
For mode amplitudes, we include terms up to 4 PN. The terms up to 3 PN are taken from [53], the 3.5 PN terms in the $(2,2)$ mode and $\ell=3$ modes are taken from [54] and [55], respectively, and the 4 PN terms in $(2,2)$ mode are taken from [47]. We also include memory terms up to 3 PN [31] for mode amplitudes.
We also generate the PN Moreschi supermomentum [56], which we will later use to fix the BMS frame of NR systems, as we will detail in Sec. II.2. The Moreschi supermomentum modes are taken from [31]. We leave the effects of eccentric orbits for discussion in Sec. VI.
To produce the PN phase as a function of time, we use the TaylorT1 approximant. The PN generation is implemented in the Python package NRPNHybridization [45].
There are 8 independent physical parameters in quasicircular PN theory: mass ratio $q$, total mass $M$, and two three-dimensional spin vectors $\bm{\chi}_{1}$ and $\bm{\chi}_{2}$. The remaining parameters are all gauge. There are 4 parameters for a time shift and $\mathrm{SO}(3)$ rotation, which were considered in previous hybridization work. The rest of the BMS group includes a Lorentz boost (3 parameters), and a supertranslation (described below), which is given by a smooth real-valued function $\alpha(\theta,\phi)$ on the sphere. We band-limit $\alpha$ to $\ell\leq 8$ modes, and its (0,0) mode is already accounted for in the time shift above, so $\alpha$ contributes an additional $(\ell+1)^{2}-1=80$ real parameters.
@@ II.2 Numerical-relativity waveforms and BMS frames
Reference [62] emphasized the importance of having both PN and NR waveforms in the same BMS frame. We follow their procedure to map NR waveforms to the corresponding PN BMS frame. First, we determine the space translation and boost by minimizing the center-of-mass charge. We then calculate the supertranslations that map the $\ell\leq 2$ modes of the NR Moreschi supermomentum $\Psi^{\mathrm{NR}}_{\mathrm{M}}$ to the corresponding PN Moreschi supermomentum $\Psi^{\mathrm{PN}}_{\mathrm{M}}$, where $\Psi_{\mathrm{M}}$ is
@@
with $\Psi_{2}$ being the Weyl scalar with spin-weight 0, $\sigma$ the Newman-Penrose shear, and an overbar representing complex conjugation. For a quantity $\eta$ of spin weight $s$, the spin-weight raising operator $\eth$ is defined as [63, 64]
@@
We use NR waveforms that are computed using a Cauchy-characteristic evolution (CCE) [40, 41], for which the Cauchy evolution parts are performed using the Spectral Einstein Code (SpEC) [65], and the CCE computation is performed using the SpECTRE code [66]. CCE uses a worldtube from the Cauchy evolution of the Einstein field equations as the inner boundary of a characteristic evolution on a null foliation. The gravitational information is propagated and evolved to future null infinity. Note that $\Psi_{2}$ and $\sigma$ required in Eq. (3) are calculated by the SpECTRE CCE code.
@@
The NR waveforms that we use are from simulations in the SXS catalog [43]. We plot the parameter space $\chi_{p}$ versus $q$ in Fig. 1, where $\chi_{p}=\text{max}\{\chi_{1\perp},\frac{4+3q}{q(4q+3)}\chi_{2\perp}\}$ [67] and $\chi_{1,2\perp}$ is the magnitude of the projection in the orbital plane. They are systems SXS:BBH:1412–1416 and 2617–2640. Of these, SXS:BBH:1412–1416 and 2617-2621 are exceptionally long, spanning over 100 orbits. This extensive length enables us to evaluate the validity of hybridization at an early stage in the binary’s inspiral. The mass ratio for both the SXS:BBH:1412–1416 and 2617-2635 systems falls within the range $101<q<$ 10. As for the SXS:BBH:2636–2640 systems, the mass ratios are approximately $q\approx 8$. The waveforms SXS:BBH:1412–1416 correspond to spin-aligned systems, while the other systems utilized in our study are precessing systems. The magnitudes of the spin parameters $\chi_{1}$ and $\chi_{2}$ are constrained as $|\chi_{1}|,|\chi_{2}|\leq 0.8$.
@@ III Matching PN and NR parameters
As we will see in Sec. V, the binary parameters used in PN and measured in NR often do not mean the same thing. Therefore, substituting NR parameters for those in the PN expansion will result in an inaccuracy that can cause the differences in the matching window between PN and NR waveforms to be one to three orders of magnitude larger. Therefore, we must identify the proper PN parameters that correspond to the NR systems.
@@
Figure 4: Strain and residual of two PN waveforms, for a test in which we match the two waveforms by varying the parameters of one of them. The target PN waveform has parameters $q=1.5,\,M=1.0,\,\bm{\chi}_{1}=(0.5,0.2,0.3),\,\bm{\chi}_{2}=(0.1,0.3,0.4)$, and is rotated away by $\bm{\mathrm{R}}=\exp(0.05,0.02,0.03)$ from the original frame in which the PN waveform is generated. The initial guess is $q=1.65,\,M=1.05,\,\bm{\chi}_{1}=(0.42,0.24,0.47),\,\bm{\chi}_{2}=(0.01,0.20,0.41)$. The matching window is the region between two vertical dotted lines, which is $5000\,M$ long.

### v2

@@ Abstract
Numerical relativity (NR) simulations of binary black holes provide precise waveforms, but are typically too computationally expensive to produce waveforms with enough orbits to cover the whole frequency band of gravitational-wave observatories. Accordingly, it is important to be able to hybridize NR waveforms with analytic, post-Newtonian (PN) waveforms, which are accurate during the early inspiral phase. We show that to build such hybrids, it is crucial to both fix the Bondi-Metzner-Sachs (BMS) frame of the NR waveforms to match that of PN theory, and optimize over the PN parameters to mitigate the error caused by the discrepancy between NR and PN parameters. We test such a hybridization procedure including all spin-weighted spherical harmonic modes with $|m|\leq\ell$ for $\ell\leq 8$, using 29 NR waveforms with mass ratios $q\leq 10$ and spin magnitudes $|\chi_{1}|,|\chi_{2}|\leq 0.8$. We find that for spin-aligned systems, the PN and NR waveforms agree very well. The difference is limited by the small nonzero orbital eccentricity of the NR waveforms, or equivalently by the lack of eccentric terms in the PN waveforms. To maintain full accuracy of the simulations, the matching window for spin-aligned systems should be at least 5 orbits long and end at least 15 orbits before merger. For precessing systems, the errors are larger than for spin-aligned cases. The errors are likely limited by the absence of precession-related spin-spin PN terms. Using $10^{5}\,M$ long NR waveforms, we find that there is no optimal choice of the matching window within this time span, because the hybridization result for precessing cases is always better if using earlier or longer matching windows. We provide the mean orbital frequency of the smallest acceptable matching window as a function of the target error between the PN and NR waveforms and the black hole spins.
@@ I Introduction
Such hybridization procedures were developed, for example, in Refs. [25, 26, 27, 28, 29, 30]. These procedures attained mismatches of $10^{-4}$ between the hybrid and NR waveforms in the matching window (for $m\neq 0$ modes) for non-precessing systems [29], and between $10^{-2}$ and $10^{-1}$ (for $m\neq 0$ modes) for precessing systems [30]. However, these procedures have four major limitations: (1) They did not account for the $m=0$ modes since their NR waveforms did not include the correct gravitational-wave memory contribution. (2) The NR waveforms used in these procedures were not in the same frame as the PN waveforms. Reference [31] showed that comparing CCE waveforms without fixing the gauge freedom can result in relative errors three orders of magnitude larger than if one first fixes such freedoms. And they also showed that comparing CCE waveforms after mapping them to the same BMS frame can achieve smaller errors than comparing the extrapolated waveforms. Furthermore, Ref. [31] successfully developed a hybrid procedure that incorporates memory-containing waveforms. (3) These procedures used NR values of masses and spins to compute the corresponding PN waveform. However, the mass and spin parameters are defined differently in PN than they are in NR, and hence the NR and PN parameters are expected to agree only in the limit of infinite binary separation, even for infinite-resolution NR and infinite-order PN. In addition, for finite-resolution NR and finite-order PN there are systematic errors in the mapping from parameters to waveforms. (4) Existing procedures are primarily applicable to spin-aligned or anti-aligned systems, such as binaries born from isolated stellar evolution [32, 33, 34]. The accuracy of hybridization for precessing systems is severely limited by the inaccurate methodology previously mentioned: the mismatch is on the order of $10^{-2}$ or higher [30]. Theoretical considerations [32, 34] and observational evidence [9, 35, 36] indicate that spins can be randomly oriented for binaries that form dynamically in dense environments, and even binaries formed in isolation can exhibit distinctive precessional dynamics. Therefore, it is crucial to develop accurate hybridization procedures that are valid for precessing systems.
@@
One technique that enables the extraction of waveforms at null infinity is Cauchy-characteristic evolution (CCE) [37, 38, 39, 40, 41]. It uses a worldtube from the Cauchy evolution of the Einstein field equations as the inner boundary to a characteristic evolution along null rays. The gravitational information is propagated and evolved to future null infinity, so one can obtain waveforms in any desired gauge. Reference [31] has shown that by mapping the CCE waveform to the correct Bondi-Metzner-Sachs (BMS) frame, the NR waveform agrees much better with the PN waveform. A key advantage of the hybrid waveforms presented here is that we use CCE NR waveforms. By contrast, previous hybrid waveform models [25, 26, 27, 28, 29, 30] used NR waveforms that were computed at null infinity using an extrapolation technique [42] that does not have a well-defined BMS frame and does not yield the correct memory contribution.
PN parameter spaceNR parameter spaceWaveform space ($\infty$-dim)$h_{\text{PN}}$$h_{\text{NR}}$efer to captionRefer to caption
Figure 1: NR and PN parameter spaces are distinct because of the different gauges and different definitions of parameters used in PN and NR models. The maps into infinite-dimensional waveform space differ at each NR resolution and PN order. We optimize PN parameters and adjust the BMS frame to minimize the “vertical” difference transverse to the images of NR and PN parameter spaces.
@@
In this work, we simultaneously optimize over the PN parameters and fix the BMS frame of NR waveforms to build hybrid waveforms, so that we can mitigate the error caused by the discrepancy between NR and PN parameters, and inaccurate BMS frame. This is depicted graphically in Fig. 1. The optimization and the fixing of the BMS frame, which includes the alignment of the rotation and time shift, are performed inside a time interval called a matching window. Because we use CCE and thus obtain the correct gravitational-wave memory, we are able to include all $-\ell\leq m\leq\ell$ for $\ell\leq 8$ spin-weighted spherical harmonic modes in the hybrid waveforms. We test the hybridization procedure using 29 NR waveforms from the SXS catalog [43] with mass ratios $q\leq 10$ and $|\chi_{1}|,|\chi_{2}|\leq 0.8$. The length of these waveforms ranges from 60 to over 100 orbits, making it possible to test the hybrid waveforms by comparing them to long NR waveforms at a very early stage. We find that for spin-aligned systems, we can reduce the PN–NR mismatches to the error caused by small non-zero eccentricities of the NR systems. We achieve a mismatch (in a 10-orbit long matching window ending $8000\,M$ before merger) around two to three orders of magnitude smaller than previous models [25, 26, 27, 28, 29]. For precessing systems, we find that hybridization is likely limited by the absence of precession-related spin-spin PN terms, yielding mismatches one to two orders of magnitude smaller than previous models [30].
@@
This paper is organized as follows. In Sec. II, we introduce the PN and NR waveforms that we use to build the hybrid waveforms, and we discuss the freedoms in the physical parameters, 3D rotations and time shift, and further BMS frame parameters not previously considered. In Sec. III, we explain why there are discrepancies between PN and NR parameters, and how we match the physical parameters and inertial frames between NR and PN waveforms. Following this, we present the complete procedure to build the hybrid waveforms in Sec. IV. In Sec. V, we show the hybridization results as well as the improvement from PN parameters and BMS frame fixing compared with previous models, and we use the findings to more precisely compare the PN and NR waveforms. Sec. VI analyzes the sources of error for spin-aligned and precessing systems. Additionally, based on our error analysis, we make recommendations for future surrogate models’ matching window selection in Sec. VII. In Sec. VIII, we provide a few closing remarks. We make the code used for this hybridization procedure publicly available in Ref. [44].
@@ II.1 Post-Newtonian waveforms
In this study, we use PN theory to describe the GW waveforms of BBHs in quasi-circular orbits during the early inspiral stage, and include GW memory contributions. Memory effects should be inherent in any correct PN formulation, but their importance can often be overlooked. In our case, we find that accounting for memory is crucial when comparing to NR waveforms which have memory (since memory effects inherently arise when directly solving Einstein’s equations), and thus crucial for accurate hybridization.
@@
For mode amplitudes, we include terms up to 4 PN. The terms up to 3 PN are taken from [52], the 3.5 PN terms in the $(2,2)$ mode and $\ell=3$ modes are taken from [53] and [54], respectively, and the 4 PN terms in $(2,2)$ mode are taken from [46]. We also include non-linear memory terms up to 3.5 PN [55, 31] for mode amplitudes, but they are only complete at 3 PN for non-spinning systems, and 2 PN for spinning systems. They are computed using the procedure outlined in Ref. [31].
We also generate the PN Moreschi supermomentum [56], which we will later use to fix the BMS frame of NR systems, as we will detail in Sec. II.2. The Moreschi supermomentum modes are taken from [31], which has the leading order spin-spin and spin-orbit terms. We leave the effects of eccentric orbits for discussion in Sec. VI.
To produce the PN phase as a function of time, we use the TaylorT1 approximant. The difference between different PN approximants is anticipated to be at the same order with the PN truncation error, which is controlled by the PN truncation order, implying that switching PN approximants is unlikely to substantially reduce the hybridization error. The PN generation is implemented in the Python package NRPNHybridization [44].
There are 8 independent physical parameters in quasicircular PN theory: mass ratio $q$, total mass $M$, and two three-dimensional spin vectors $\bm{\chi}_{1}$ and $\bm{\chi}_{2}$. The remaining parameters are all gauge. There are 4 parameters for a time shift and $\mathrm{SO}(3)$ rotation, which were considered in previous hybridization work. We will determine the above 12 parameters by performing a 12D optimization. The rest of the BMS group includes a Lorentz boost (3 parameters), and a supertranslation (described below), which is given by a smooth real-valued function $\alpha(\theta,\phi)$ on the sphere. We band-limit $\alpha$ to $\ell\leq 8$ modes, and its (0,0) mode is already accounted for in the time shift above, so $\alpha$ contributes an additional $(\ell+1)^{2}-1=80$ real parameters. We will discuss how to determine these 80 parameters in the next subsection.
@@ II.2 Numerical-relativity waveforms and BMS frames
Note that the BMS frame is defined at null infinity, but the outer boundary of NR simulations is set at a finite radius, which doesn’t have the information from null infinity. So the NR BMS frame can not be specified arbitrarily when setting up the simulation, and therefore in practice an NR simulation (before any BMS frame fixing is done) is in some arbitrary uncontrolled BMS frame. Reference [62] emphasized the importance of having both PN and NR waveforms in the same BMS frame. We follow their procedure to map NR waveforms to the corresponding PN BMS frame. First, we determine the space translation and boost by minimizing the center-of-mass charge. We then calculate the supertranslations that map the $\ell\geq 2$ modes of the NR Moreschi supermomentum $\Psi^{\mathrm{NR}}_{\mathrm{M}}$ to the corresponding PN Moreschi supermomentum $\Psi^{\mathrm{PN}}_{\mathrm{M}}$, where $\Psi_{\mathrm{M}}$ is
@@
with $\Psi_{2}$ being the Weyl scalar with spin-weight 0, $\sigma$ the Newman-Penrose shear, and an overbar representing complex conjugation. For a quantity $\eta$ of spin weight $s$, the spin-weight raising operator $\eth$ can be written as [63, 64]
@@
We use NR waveforms that are computed using a Cauchy-characteristic evolution (CCE) [40, 41], for which the Cauchy evolution parts are performed using the Spectral Einstein Code (SpEC) [65], and the CCE computation is performed using the SpECTRE code [66]. CCE uses a worldtube from the Cauchy evolution of the Einstein field equations as the inner boundary of a characteristic evolution on a null foliation. The gravitational information is propagated and evolved to future null infinity. Note that $\Psi_{2}$ and $\sigma$ required in Eq. (3) are calculated by the SpECTRE CCE code. Note that unlike CCE waveforms, extrapolated waveforms obtained using the methods described in Ref. [42] cannot be mapped to the PN BMS frame, because they do not have an accurate value of $\Psi_{2}$ for calculating the Moreschi supermomentum.
@@
The NR waveforms that we use are from simulations in the SXS catalog [43, 67]. We plot the parameter space $\chi_{p}$ versus $q$ in Fig. 2, where $\chi_{p}=\text{max}\{\chi_{1\perp},\frac{4+3q}{q(4q+3)}\chi_{2\perp}\}$ [68] and $\chi_{1,2\perp}$ is the magnitude of the projection in the orbital plane. They are systems SXS:BBH:1412–1416 and 2617–2640. Of these, SXS:BBH:1412–1416 and 2617-2621 are exceptionally long, spanning over 100 orbits. This extensive length enables us to evaluate the validity of hybridization at an early stage in the binary’s inspiral. The mass ratio for both the SXS:BBH:1412–1416 and 2617-2635 systems falls within the range $101<q<$ 10. As for the SXS:BBH:2636–2640 systems, the mass ratios are approximately $q\approx 8$. The waveforms SXS:BBH:1412–1416 correspond to spin-aligned systems, while the other systems utilized in our study are precessing systems. The magnitudes of the spin parameters $\chi_{1}$ and $\chi_{2}$ are constrained as $|\chi_{1}|,|\chi_{2}|\leq 0.8$.
@@ III Matching PN and NR parameters
There are three reasons why we want to perform the optimization over all intrinsic parameters:
  1. 1.
    The definitions of PN parameters are not consistent with the NR definitions. The NR parameters are measured quasi-locally on the apparent horizon, as detailed in [67]. Meanwhile PN parameters are attributed to point particles, via asymptotic matching in “body zones”, as explained in [69]. The difference between the PN and NR definitions should be small as we see in the optimization results, but perfect agreement between them is anticipated only at infinite separation, even for infinite-order PN and infinite-resolution NR. The leading order contribution to the difference is expected to be the binding energy between the two black holes, which appears at 1 PN order. Determining the exact relationship between PN and NR definitions requires solving for the apparent horizon in the PN metric and evaluating the parameters at the PN apparent horizon, which is beyond the scope of this paper. We are working on this issue as a separate project.
    The difference in parameter definitions is not an error or uncertainty, but a conventional discrepancy, thus it has no physical significance. For astrophysical parameter inference, it is recommended to quote PN parameters rather than NR, since these correspond to masses and spins infinitely far in the past.
  2. 2.
    Even if we used the same definitions of mass and spin quantities in PN and NR, the definitions are gauge dependent. Apparent horizon integrals used in NR are gauge dependent since apparent horizons are themselves slicing-dependent [70]. Meanwhile specifying a spin in PN requires a choice of spin supplementary condition (SSC), which is a type of gauge choice [71]. A choice of SSC also affects the definition of mass in PN. This gauge-dependence also applies to the total binary mass, which is distinct from the gauge-invariant ADM energy; these differ by the binding energy and the energy carried away by radiation.
  3. 3.
    Both PN and NR waveforms have systematic errors. The PN errors are mainly due to the truncation at a finite PN order, and the NR errors are mainly due to numerical truncation. The error is supposed to affect both the waveform and the parameters. As we will show in Sec. V.1, the difference between the PN and NR parameters are at the same order as the residual between the strain modes. Therefore, there is no reason to force the difference between the PN and NR parameters to be zero, seeing as this would introduce more error in the strain and thus make the hybrid waveform less accurate. Given the finite length and finite resolution of the waveforms, the optimized parameters can compensate for the errors in the model and the unmodeled effects, to provide the best agreement and smoothest transition between the NR and PN waveforms.
Therefore, we must identify the proper PN parameters that correspond to the NR systems.
@@
Figure 5: Strain and residual of two PN waveforms, for a test in which we match the two waveforms by varying the parameters of one of them. The target PN waveform has parameters $q=1.5,\,M=1.0,\,\bm{\chi}_{1}=(0.5,0.2,0.3),\,\bm{\chi}_{2}=(0.1,0.3,0.4)$, and is rotated away by $\bm{\mathrm{R}}=\exp(0.05,0.02,0.03)$ from the original frame in which the PN waveform is generated. The initial guess is $q=1.65,\,M=1.05,\,\bm{\chi}_{1}=(0.42,0.24,0.47),\,\bm{\chi}_{2}=(0.01,0.20,0.41)$. The parameters obtained by the optimization are $q=1.500004,\,M=1.000005,\,\bm{\chi}_{1}=(0.10001,0.200002,0.30003),\,\bm{\chi}_{2}=(0.399994,0.09998,0.20006)$. The matching window is the region between two vertical dotted lines, which is $5000\,M$ long.